At 0.0°C and 1.00 atm, 1.00 mol of a gas occupies a volume of 0.0224 m3.

If the gas is nitrogen (N2), the principal component of air, what is the mass density? Molar mass of Nitrogen is 14.00674 g/mol.


1
Expert's answer
2022-05-18T15:28:15-0400

"pV=\frac mM RT,\implies"

"m=\frac{MpV}{RT}=14.01~g."

"\rho=\frac mV=0.625~\frac{kg}{m^3}."